OpenStudy (supercait10):

Can you help?? ( Please read: Have to put the questions and choices down in the comments)

OpenStudy (danjs):

i'll try

OpenStudy (supercait10):

Smiplify: \[(^{y4} ) \] \[(^{y3} )\] A. \[(^{y1})\] B. \[(^{y7})\] C. \[(^{y12})\] D. \[(^{y43})\]

OpenStudy (danjs):

sorry, just got back

OpenStudy (danjs):

what is the operation y^4 and y^3

OpenStudy (supercait10):

yup

OpenStudy (danjs):

a fraction?

OpenStudy (danjs):

\[\frac{ y^4 }{ y^3 }\] ?

OpenStudy (supercait10):

no just simplify y^4 y^3

OpenStudy (danjs):

\[y^4y^3\] that must be multiplication then y^4 * y^3

OpenStudy (danjs):

From the exponents of like bases rules, if you multiply the terms, you add together the powers to combine

OpenStudy (supercait10):

idk all it says is Simplify: (y^4) (y^3) that all it says

OpenStudy (danjs):

\[y^4 *y^3 = y ^{4+3} = y^7\]

OpenStudy (supercait10):

okay then can u help me on the rest

OpenStudy (danjs):

You just have to remember the 4 or so rules for combining exponents

OpenStudy (supercait10):

okay

OpenStudy (danjs):

a^b * a^c = a^(b+c)

OpenStudy (danjs):

sure post another if you want

OpenStudy (danjs):

if you trying to type all those and dont like it, you can always hit the draw button, and just do that

OpenStudy (supercait10):

sorry my laptop is being stupid